# Explainable Machine Learning Identification of Superconductivity from Single-Particle Spectral Functions

### v1

@@ Abstract
Abstract: The traditional method of identifying symmetry-breaking phase transitions through the emergence of a single-particle gap encounters significant challenges in quantum materials with strong fluctuations. To address this, we have developed a data-driven approach using a domain-adversarial neural network trained on simulated spectra of cuprates. This model compensates for the scarcity of experimental data – a significant barrier to the wide deployment of machine learning in physical research – by leveraging the abundance of theoretically simulated data. When applied to unlabeled experimental spectra, our model successfully distinguishes the true superconducting states from gapped fluctuating states, without the need for fine temperature sampling across the transition. Further, the explanation of our machine learning model reveals the crucial role of the Fermi-surface spectral intensity even in gapped states. It paves the way for robust and direct spectroscopic identification of fluctuating orders, particularly in low-dimensional, strongly correlated materials.
@@
Understanding and controlling the materials’ physical properties are key pursuits of the quantum materials research today. Since the electronic structure of materials underpins many physical properties, the single-particle spectral function proves to be an effective quantity for their characterization [1]. This function captures the probability of an individual electron occupying a specific energy-momentum state in a many-electron system. While it cannot replace the full many-body wavefunction, it provides insights into the low-energy properties such as conductivity and thermal excitations. Notably, the single-particle spectral function can be directly measured using angle-resolved photoemission spectroscopy (ARPES). Both lab-based and synchrotron-based ARPES techniques have been extensively employed to investigate emergent electronic states in materials [2, 3], substantially accelerating the advancement in the field of quantum materials [4].
@@
Figure 1: Schematic illustrating the single-particle gap in materials experiencing strong fluctuations. At low temperatures (red), the material exhibits a nonzero order parameter $\langle\Delta\rangle\neq 0$ and a well-defined single-particle gap. As the temperature increases above $T_{c}$ (green), the local excitations lose long-range coherence and the average order parameter $\langle\Delta\rangle=0$. However, the system still displays pronounced fluctuations with a non-negligible $\langle\Delta^{2}\rangle$, leading to a finite single-particle gap. These short-range fluctuations gradually diminish with further temperature increase, ultimately resulting in a normal state (blue).
@@
Figure 2: Model architecture. Left to right: Consecutive convolutional stacks as the feature extractor convert the ARPES spectrum (using experimental data for BSCCO OD58 as an example) into feature maps with various convolutional kernels. A pooling layer and activation are then applied to compress the feature maps and pass the data to the next layer. After four convolutional layers, the feature maps are pooled and flattened, then directed into two distinct fully connected neural networks: The label predictor (blue) classifies the spectra into the superconducting (SC) and normal phases; The domain classifier (pink) classifies the sample source either from simulation or experiment. It connects to the feature extractor via a gradient reversal layer that multiplies the gradient by a certain negative constant during the backpropagation (dashed line). The number and size of each layer plotted in this figure are for illustrative purposes.
As investigations into quantum materials deepen, new challenges arise due to the inherent limitations of representing a many-body state with only its single-particle excitations. For traditional metals and semiconductors, this approximation is effective as interactions are sufficiently screened. In such scenarios, ARPES can accurately identify electronic phase transitions through a single-particle gap, stemming from the development of a (mean-field) order parameter. By measuring the energy gap size, one can deduce the strength of order parameters. However, in quantum materials, emergent phases are substantially influenced, or even governed, by entanglement among multiple particles. Consequently, these phases often deviate from traditional mean-field theory predictions. For example, transitions between phases with distinct symmetries often lead to a regime characterized by intense short-range fluctuations. The effects of these fluctuations on the single-particle spectral function are similar to that of a broken-symmetry long-range order, often reflected in the form of partial spectral weight depletion that precedes the establishment of a full energy gap (see Fig. 1). This phenomenon has been widely observed in Mott insulators, charge density wave systems, and unconventional superconductors above their thermodynamic phase transition temperatures [5, 6, 7, 8, 9]. This often renders spectroscopic probes incapable of detecting thermodynamic phase transitions, despite the rapidly growing demand for in situ and in operando detection of novel electronic phases [10, 11].
Superconductivity stands out as a central topic in condensed matter research, characterized by dissipationless transport due to the formation of global phase coherence among Cooper pairs of electrons. In the mean-field framework, superconductivity is linked to a single-particle gap twice the size of the order parameter. In scenarios where direct transport measurements are impractical – such as with thin films, functionalized surfaces, nonequilibrium systems, or extreme conditions precluding direct contacts – the presence of this gap serves as a fingerprint of superconductivity [12, 13, 14]. However, in correlated materials like cuprates and FeSe, quantum fluctuations disrupt the direct link between the gap and superconductivity [5, 6, 7, 8]. In these cases, the gap-opening temperature $T_{\rm gap}$ can significantly exceed the superconductivity transition temperature $T_{c}$. There is a pressing need for an efficient approach to directly identify superconductivity from the single-particle spectra alone, to facilitate high-throughput discovery of superconductive materials. Recent studies have revealed the potential for ARPES to also encode information on superconducting phase coherence [8, 15, 16], but pinpointing this information remains challenging without tracking the full temperature evolution of quasiparticle intensity. This challenge likely stems from the oversimplification of the highly complex spectra of a many-body system.
To address this issue, we develop a machine learning (ML) model to directly classify ordered superconducting phases from ARPES spectra without needing complete temperature-dependent trends. One significant challenge we face is the limited experimental training data, a ubiquitous obstacle in artificial intelligence (AI) for science community [17]. While most state-of-the-art ML algorithms are data demanding, scientific data such as ARPES spectra are often scarce due to the complexity of experiments and the lack of standardization in data curation. A promising solution is to utilize training data curated in a simulated environment to enable efficient exploration of the parameter space in silico[18]. Inspired by this idea, we train a convolutional neural network (CNN) model using simulated ARPES spectra and then apply it to classify experimental ARPES spectra. CNN is chosen here due to its wide applications in image classification, where filters in convolutional layers distill specific patterns from two-dimensional data. However, the idealized conditions assumed in simulations often fail to replicate the experimental noise, resolutions, and laboratory condition variability, leading to poor performance when applying simulation-trained models to actual experimental data. To overcome this experiment-simulation discrepancy, we employ domain adaptation techniques to adversarially train the CNN.
As illustrated in Fig. 2, the domain adversarial neural network (DANN) [19] adapted in this work comprises a feature extractor, a domain classifier, and a label predictor. The feature extractor extracts a shared feature representation of both simulated and experimental spectra. The domain classifier then determines whether the extracted features originate from the simulated or experimental datasets. The label predictor utilizes these features and outputs the corresponding material sample’s phase label, which can be either superconducting (SC) or normal phase. Unlike traditional CNN, which comprises only a feature extractor and a label predictor, the domain classifier discourages the feature extractor from generating distinguishable representations between simulated and experimental data. This setup promotes the emergence of latent representations that are shared by both simulated and experimental spectra during optimization. Importantly, the label predictor is trained exclusively on the labeled simulation data (SC vs normal), with no phase information required for the experimental spectra. Thus, the model is transferable to classify experimental spectra at any random temperature. To enable phase label prediction, a softmax layer is added to the label predictor to convert the predicted logits $\mu_{\alpha}$ for each phase $\alpha$ ($\alpha$ = SC or normal) into probability values $p_{\alpha}\propto e^{\mu_{\alpha}}$, obeying the sum rule $\sum_{\alpha}p_{\alpha}=1$.
The experimental ARPES spectra were collected for two cuprate samples: a super-oxygenated Bi2Sr2CaCu2O8 (BSCCO) with $T_{c}=50$ K (OD50) and an oxygenated BSCCO with $T_{c}=58$ K (OD58). A total of 41 and 45 spectra are measured in these two samples, respectively, spanning over a wide temperature range from 14 K to 102 K and both superconducting and normal phases (reused from Refs. 20 and 8). To better distinguish fluctuations encoded in spectral features, we applied a uniform background removal approach to both simulated and experimental data (see Supplementary Note 1). Fig. 3a presents four spectra taken from the OD50 sample after the background removal. While superconductivity, determined by transport measurements, disappears when the temperature exceeds 50 K, the single-particle gap remains open until $T_{\rm gap}=65$ K [20]. It is important to note that the true label, or the phase corresponding to each experimental spectrum, is invisible to our ML model. All labeled training data are obtained from simulations following the method outlined in Ref. 8 (see Methods for details).
@@
Figure 3: Classification of superconducting phase using the DANN model. a Experimental ARPES spectra for the BSCCO OD50 sample at four different temperatures, with dashed lines denoting the Fermi level. The upper and lower bars indicate the phase labels, independently determined by experiments and invisible to the ML model, and the gap sizes. b,c Confusion matrices obtained for binary classification of the ARPES spectra collected from b BSCCO OD50 and c OD58 samples at various temperatures, yielding accuracies of 97.6% and 97.8%, respectively. d,e The ML-predicted SC probability $\bar{p}_{\textrm{SC}}$ for spectra obtained from the d BSCCO OD50 and e OD58 samples, respectively. The $\bar{p}_{\textrm{SC}}$s are calculated for each spectrum but are sorted here by their experimental temperatures, which are unknown to the ML model. The classification is based on whether $\bar{p}_{\textrm{SC}}$ exceeds 0.5. Correctly classified spectra are denoted by solid red and blue dots, while misclassified data are depicted by open dots. The red and blue lines indicate experimentally determined transition temperature $T_{c}$ and gap-opening temperature $T_{\rm gap}$ (both invisible to the ML model), respectively.
@@
Figures 3b and c show the phase classification results for the BSCCO OD50 and OD58 experimental spectra. The model’s predictions are represented by a scalar “SC probability” $\bar{p}_{\textrm{SC}}$ for each input spectrum. Such an SC probability $\bar{p}_{\textrm{SC}}$ is obtained from an ensemble approach, where $\bar{p}_{\textrm{SC}}$ for each spectrum is the averaged $p_{\textrm{SC}}$ over 10 different DANN models with various initializations. Normalized by the sum rule, a sample corresponding to an input spectrum is predicted as SC when $\bar{p}_{\textrm{SC}}$ exceeds 50%. For the BSCCO OD50 sample, only one spectrum measured in the normal phase is misclassified as SC. Conversely, all other 40 experimental spectra are correctly classified, yielding an accuracy of 97.6%. The performance is similar for the OD58 sample, where only one spectrum in the SC state is misclassified.
@@
As the experimental spectra are obtained for two material samples at different temperatures, we further analyze the predicted SC probability $\bar{p}_{\textrm{SC}}$ as a function of temperature in Figs. 3c and d. Although our ML model independently classifies each ARPES spectrum and does not have access to temperature information, we observe a roughly monotonic relationship between $\bar{p}_{\textrm{SC}}$ and the actual temperature $T$. The overall shape of the $\bar{p}_{\textrm{SC}}(T)$ resembles an inverted sigmoid function: it shows constantly high values (close to 1) at low T and low values (close to 0) at high T, with a rapid transition from 1 to 0 near the experimentally determined $T_{c}$. The only misclassified data point in Fig. 3c corresponds to a sample measured at a temperature close to $T_{c}$. Given that none of the experimental spectra are labeled in ML, the monotonic decrease in the predicted probability with increasing temperature indicates that our model has successfully captured the superconducting phase ordering encoded in the details of ARPES spectra. Interestingly, the predicted $\bar{p}_{\textrm{SC}}$ does not exhibit any anomaly near the $T_{\rm gap}$. These observations reflect that the gap opening is not the sole or even primary indicator of phase transition in quantum materials. The spectral details beyond the mere presence of a gap contain extensive information about the underlying phases.
@@
It is worth noting that the simulated and experimental ARPES spectra exhibit fundamental differences, as shown in Fig. 2, due to the simplicity of the single-band model. Thus, the domain adaptation technique is crucial in bridging the gap between the simulated training spectra and experimental test spectra, hence ensuring our ML phase classifier’s high accuracy. Without the adversarial training enabled by the domain classifier, the CNN model yields a representation space where the simulated and experimental spectra occupy distinct regions of the latent space, as visualized by the t-distributed stochastic neighbor embedding [21] (t-SNE) in Fig. 4a. This separation reflects the intrinsic differences between simulated and experimental data. As a result, the classification rules learned from the simulated spectra are not directly applicable to experimental spectra. In contrast, the domain classifier in DANN significantly enhances the alignment between simulated and experimental spectra in the representation space (see Fig. 4b), providing transferable classification rules between simulated and experimental spectra. Consequently, the classification results show a substantial improvement by 19.0% and 19.2% in average accuracy on the BSCCO OD50 and OD58 samples, respectively (see Supplementary Note 2).
@@
Figure 5: Saliency distribution for experimental spectra. a A sample ARPES spectrum obtained from BSCCO OD50 at 38 K (superconducting phase). b The saliency values across various binding energies and temperatures for the BSCCO OD50 sample, quantifying the sensitivity of the ML prediction for each energy. c An ARPES spectrum from BSCCO OD58 at 34 K (superconducting phase). d Saliency values for the BSCCO OD58 sample. Grey dashed lines denote the Fermi level, while red and blue lines indicate the $T_{c}$ and $T_{\rm gap}$, as independently determined by experiments.

### v2

# Detecting Thermodynamic Phase Transition via Explainable Machine Learning of Photoemission Spectroscopy

@@ Abstract
Abstract: Identifying thermodynamic signatures of electronic phases, such as superconductivity, is challenging in low-dimensional materials due to strong fluctuations and low probing volume. Spectroscopic methods are often used to identify new bulk phases, but their main measurable quantity – electronic energy gaps – is no longer an effective order parameter in low dimensional and fluctuating systems. Combining angle-resolved photoemission with a domain-adversarial neural network, we report a data-driven method to identify thermodynamic phase transitions solely based on single-particle spectra. We demonstrate 97.6% accuracy in cuprate superconductor Bi2Sr2CaCu2O8+δ with strong superconducting fluctuations. This model notably compensates for the scarcity of experimental data by leveraging virtually inexhaustible simulated data. Further, its explainability reveals the crucial role of in-gap spectral weight in detecting phase fluctuations and thermodynamic transitions. Our work pinpoints the spectroscopic signatures of fluctuating orders and enables using spectroscopy for machine-learning-assisted material discovery for low-dimensional and strong coupling systems.
@@
I Introduction
@@
Figure 1: Schematic illustration of thermodynamic phase transition with pronounced fluctuations. A At low temperatures (red), the material exhibits a nonzero order parameter $\langle\Delta\rangle\neq 0$ and a well-defined single-particle gap. As the temperature increases above $T_{c}$ (green), the local excitations lose long-range coherence and the average order parameter $\langle\Delta\rangle=0$. However, the system still displays pronounced fluctuations with a non-negligible $\langle\Delta^{2}\rangle$, leading to a finite single-particle gap. These short-range fluctuations gradually diminish with further temperature increase, ultimately resulting in a normal state (blue). B The domain discrepancy between the scarce, noisy experimental data and the abundant, idealized simulated spectra.
New states of matter are typically identified by the emergence of long-range order via a thermodynamic phase transition. As the system crosses the critical temperature $T_{c}$, singularities in properties like correlation length and specific heat clearly delineate the transition between the ordered (broken-symmetry) phase and the normal (high-symmetry) phase. Traditional thermodynamic probes, such as calorimetry and dilatometry, rely on bulk material measurements and generally lack spatial resolution, making them unsuitable for the characterization of thin van der Waals materials or wafer-scale epitaxial materials. In contrast, electron or optical spectroscopy can be employed to measure the energy gap $\Delta$ (Fig. 1A), which serves as an indicator of long-range order under the mean-field approximation [1]. Unlike thermodynamic probes, spectroscopy offers the advantages of being non-contact, compatible with in situ and operando measurements, and is able to resolve down to micron-level spatial precision in a high-throughput manner [2, 3]. However, many low-dimensional and correlated materials exhibit significant quantum fluctuations in the normal state [4, 5, 6, 7]. Here, the global symmetry is preserved ($\langle\Delta\rangle=0$), but the spectral gap already opens ($\langle\Delta^{2}\rangle\neq 0$) which severely undermines the effectiveness of spectroscopy in distinguishing phases [8, 9]. In these situations, traditional spectroscopic approach obtains the gap-opening temperature (denoted as $T_{\rm gap}$), which may deviate substantially from the actual thermodynamic transition [10, 11, 4, 12, 13, 5, 7].
Superconductivity (SC) in low-dimensional materials serves as a prime example of how fluctuations can complicate the identification of thermodynamic transitions through spectroscopy. In a well-defined SC phase, global phase coherence among Cooper pairs eliminates electrical resistance, establishing long-range order. Under mean-field approximations, the emergence of the SC phase is linked to a single-particle gap that is generally twice the size of the order parameter. Thus, in scenarios where direct transport measurements are not feasible – such as in ultrathin films, functionalized surfaces, nonequilibrium systems, or extreme conditions – the presence of this gap often serves as a fingerprint of SC. [14, 10, 15]. However, in correlated materials like cuprates and monolayer FeSe, quantum and thermal fluctuations disrupt the straightforward relationship between the energy gap and the SC phase, causing a pronounced separation between $T_{\rm gap}$ and $T_{c}$[4, 12, 13, 5]. Intriguingly, recent studies suggest that angle-resolved photoemission spectroscopy (ARPES) could provide insights into SC phase coherence [5, 6, 16], implying that electronic spectra encode many-body information far beyond mean-field descriptions. However, extracting this information remains extremely challenging, requiring material-specific experimental setups, extensive measurements, and detailed temperature-dependent analysis of quasiparticle spectra. The difficulty reflects the pressing need to develop a general method capable of reliably probing thermodynamic phase transitions and the spectral signatures of fluctuating orders with limited experimental data.
Here, we develop a machine learning (ML) model to directly classify actual thermodynamic phase transitions from individual ARPES spectra without needing extensive experimental data or temperature-dependent analysis. One of the core obstacles in applying artificial intelligence (AI) to materials science is the scarcity of labeled experimental data, a ubiquitous issue across scientific fields [17]. While most modern ML techniques rely heavily on large datasets, ARPES spectra are particularly challenging to acquire in sufficient quantities due to the inherent complexity of the experiments and the lack of standardized data curation processes. One potential solution is to generate training data through simulations, enabling efficient exploration of parameter space in silico[18]. However, simulated data, even when augmented with synthetic noise, fall short of replicating real-world experimental noise, resolution constraints, and lab-specific variability (see Fig. 1B), leading to poor performance when applying simulation-trained models to actual experimental data. To address this experiment-simulation discrepancy, we employ an adversarial training strategy to ensure that the ML model learns classification rules transferable between simulated and experimental ARPES spectra. By leveraging this approach, our model accurately classifies SC states without needing labeled experimental data during training. Furthermore, occlusion-based attribution analysis reveals that the spectral weight distribution near the Fermi level, rather than simply detecting the presence of a gap, is key in distinguishing between SC and normal phase. This approach provides a pathway for integrating simulated single-particle functions with experimental ARPES spectra via ML, with broader applications in the study of thermodynamic phase transitions in quantum materials where labeled experimental data are scarce.
II Results
@@
Figure 2: Model architecture. Left to right: Consecutive convolutional stacks as the feature extractor convert the ARPES spectrum (using experimental data for BSCCO OD58 as an example) into feature maps with various convolutional kernels. A pooling layer and activation are then applied to compress the feature maps and pass the data to the next layer. After four convolutional layers, the feature maps are pooled and flattened and then directed into two distinct fully connected neural networks: the phase classifier (blue) classifies the spectra into the superconducting (SC) and normal phases; The domain classifier (pink) classifies the sample source either from simulation or experiment. It connects to the feature extractor via a gradient reversal layer that multiplies the gradient by a certain negative constant during the backpropagation (dashed line). The number and size of each layer plotted in this figure are for illustrative purposes. The color scale bar (bottom left) indicates the spectral intensity of the input ARPES spectra, ranging from minimum (blue) to maximum (white).
II.1 Model setup and ARPES data curation
A typical ARPES spectrum collected along a fixed momentum cut is shown in Fig. 1B. The spectral weight distribution along the horizontal and vertical axes reflects the momentum and energy of a single electron inside the material. Convolutional neural networks (CNNs) are specifically suited to analyze the ARPES spectra because the convolution operation can preserve the two-dimensional (2D) structure of momentum-energy information, instead of flattening the spectrum into a vector (see Note S1 and Figure S1 for details). However, standard CNN models’ performance can be significantly hampered by the “domain shift” issue in our case – the training data (simulated spectra) and test data (experimental spectra) are from different sources. Hence, we need to use a modified CNN version, the domain-adversarial neural network (DANN) [19], to facilitate cross-domain learning. In this work, as illustrated in Fig. 2, we adapt a DANN architecture that consists of three main components: a feature extractor, a domain classifier, and a phase classifier.
The feature extractor, implemented through convolutional layers, extracts feature representations from both simulated and experimental ARPES spectra. These representations are then passed to the phase classifier, which predicts the material phase as either SC or normal. During training, the model parameters of the feature extractor and the phase classifier are optimized by minimizing the phase label classification loss. The additional domain classifier addresses the domain shift issue through an adversarial training process, encouraging the network to discover domain-invariant latent representations shared between simulated and experimental spectra. Specifically, the domain classifier determines whether the extracted features originate from the simulated or experimental data via minimizing the cross-entropy loss, while the feature extractor is trained to confuse the domain classifier (see domain adaptation in the methods). Upon completion of the DANN training, the model is able to classify any single simulated or experimental ARPES snapshot.
Notably, to work with limited experimental data, the supervised learning of the phase classifier is entirely guided by simulated spectra with phase labels. However, unlabeled experimental spectra are used to train the domain classifier to ensure that a robust latent representation is learned by the feature extractor. Since the adversarial training process ensures domain transferability, the predicted phase labels for experimental data are expected to reflect the true material phases, even without direct supervision during training.
@@ II.1 Model setup and ARPES data curation
Figure 3: Classification of superconducting phase using the DANN model. A Experimental ARPES spectra for the BSCCO OD50 sample at four different temperatures, with dashed lines denoting the Fermi level. The upper and lower bars indicate the phase labels, independently determined by experiments and invisible to the ML model, and the gap sizes. The color scale bar indicates spectral intensity, ranging from minimum (blue) to maximum (white).B,C Confusion matrices obtained for binary classification of the ARPES spectra collected from B BSCCO OD50 and C OD58 samples at various temperatures, yielding accuracies of 97.6% and 97.8%, respectively. D,E The ML-predicted SC probability $\bar{p}_{\textrm{SC}}$ for spectra obtained from the D BSCCO OD50 and E OD58 samples, respectively. The $\bar{p}_{\textrm{SC}}$s are calculated for each spectrum but are sorted here by their experimental temperatures, which are unknown to the ML model. The classification is based on whether $\bar{p}_{\textrm{SC}}$ exceeds 0.5. Correctly classified spectra are denoted by solid red and blue dots, while misclassified data are depicted by open dots. The red and blue lines indicate experimentally determined transition temperature $T_{c}$ and gap-opening temperature $T_{\rm gap}$ (both invisible to the ML model), respectively.
The experimental ARPES spectra were collected from two cuprate samples: a super-oxygen-ated Bi2Sr2CaCu2O8 (BSCCO) with $T_{c}=50$ K (OD50) and an oxygenated BSCCO with $T_{c}=58$ K (OD58) (see details of experimental determination of $T_{c}$ in Note S6). A total of 41 and 45 spectra are measured in these two samples, respectively, spanning a temperature range from 14 K to 102 K, covering both SC and normal phases (reused from He et al. [5, 20]). Superconducting gap extraction with traditional spectral fitting method is conducted to benchmark the performance and advantage of the new method here (see Note S5 and Figure S7, and S8). All labeled training data are generated through simulations following the method outlined in simulated ARPES data. A uniform background removal process is applied to both the simulated and experimental datasets (see supplemental methods and Figure S5 for details). Fig. 3A presents four spectra taken from the OD50 sample after the background removal. Notably, while superconductivity disappears when the temperature exceeds $T_{c}=50$ K, the single-particle gap remains open until $T_{\rm gap}\sim 65$ K [20].
II.2 Superconducting phase classification
@@
The trained DANN model is applied to classify superconducting phases in BSCCO OD50 and OD58 experimental spectra, where the pronounced fluctuating gap long hinders the spectral characterization of superconducting phase with traditional spectroscopy methods [5]. As shown in Figs. 3B and 3C, the model’s predictions are represented by a scalar “SC probability” $\bar{p}_{\textrm{SC}}$ for each input spectrum. Such an SC probability $\bar{p}_{\textrm{SC}}$ is obtained from an ensemble approach, where $\bar{p}_{\textrm{SC}}$ for each spectrum is the averaged $p_{\textrm{SC}}$ over 10 different DANN models with various initializations. Normalized by the sum rule, a sample corresponding to an input spectrum is predicted as SC when $\bar{p}_{\textrm{SC}}$ exceeds 50%. For the BSCCO OD50 sample, only one spectrum measured in the normal phase is misclassified as SC. Conversely, all other 40 experimental spectra are correctly classified, yielding an accuracy of 97.6%. The performance is similar for the OD58 sample, where only one spectrum in the SC state is misclassified.
@@
As the experimental spectra are obtained for two material samples at different temperatures, we further analyze the predicted SC probability $\bar{p}_{\textrm{SC}}$ as a function of temperature in Figs. 3D and 3E. Although our ML model independently classifies each ARPES spectrum and does not have access to temperature information, we observe a roughly monotonic relationship between $\bar{p}_{\textrm{SC}}$ and the actual temperature $T$. The overall shape of the $\bar{p}_{\textrm{SC}}(T)$ resembles an inverted sigmoid function: it shows constantly high values (close to 1) at low temperatures and low values (close to 0) at high temperatures, with a rapid transition from 1 to 0 near the experimentally determined $T_{c}$. The only misclassified data point in each of Figs. 3D and 3E corresponds to a sample measured at a temperature close to $T_{c}$. Given that none of the experimental spectra are labeled in ML, the monotonic decrease in the predicted probability with increasing temperature indicates that our model has successfully captured the thermodynamic phase transition encoded in the details of ARPES spectra. Interestingly, the predicted $\bar{p}_{\textrm{SC}}$ does not exhibit any anomaly near the $T_{\rm gap}$. This is in stark contrast to traditional superconducting spectral gap fitting results, which is instead only sensitive to $T_{\rm gap}$, and furthermore suffers a rapidly exacerbating stability issue above $T_{c}$ (see Note S5 and Figure S7-S8). These observations reflect that the gap opening is not the sole or even primary indicator of a thermodynamic phase transition in many quantum materials. Instead, spectral distributions beyond the band dispersion contain extensive information about the emergence of long-range order associated with this phase transition.
@@
It is worth noting that the simulated and experimental ARPES spectra exhibit fundamental differences, as shown in Fig. 1B, due to the simplicity of the single-band model and unreplicable experimental noise. Thus, the domain adaptation technique is crucial in bridging the gap between the simulated training spectra and experimental test spectra, hence ensuring our ML phase classifier’s high accuracy. Without the adversarial training enabled by the domain classifier, the CNN model yields a representation space where the simulated and experimental spectra occupy distinct regions of the latent space, as visualized by the t-distributed stochastic neighbor embedding [21] (t-SNE) in Fig. 4A. This separation reflects the intrinsic differences between simulated and experimental data in this representation. As a result, the phase classification rules learned from the simulated spectra are not directly applicable to experimental spectra. In contrast, the domain classifier in DANN ensures a better representation is learned to capture the underlying physics rather than the apparent differences, significantly enhancing the alignment between simulated and experimental spectra in the representation space (see Fig. 4B). Consequently, the DANN provides transferable phase classification rules between simulated and experimental spectra, and the classification results show a substantial improvement by 19.0% and 19.2% in average accuracy on the BSCCO OD50 and OD58 samples, respectively (see Note S2 and Tables S1 and S2).
@@
Figure 5: Saliency distribution for experimental spectra. A A sample ARPES spectrum obtained from BSCCO OD50 at 38 K (superconducting phase). The color scale bar indicates the spectral intensity, ranging from minimum (blue) to maximum (white). B The saliency values across various binding energies and temperatures for the BSCCO OD50 sample, quantifying the sensitivity of the ML prediction for each energy. The color scale bar indicates positive saliency, from low (white) to high (black). C An ARPES spectrum from BSCCO OD58 at 34 K (superconducting phase). The color scale bar represents the spectral intensity, as in (A). D Saliency values for the BSCCO OD58 sample. The color scale is the same as in (B). Grey dashed lines denote the Fermi level, while red and blue lines indicate the $T_{c}$ and $T_{\rm gap}$, as independently determined by experiments.
II.3 Physical explanation of the ML model